# NEST: Self-supervised Fast Conformer as All-purpose Seasoning to Speech Processing Tasks

### v1

@@ Abstract
Self-supervised learning has been proved to benefit a wide range of speech processing tasks, such as speech recognition/translation, speaker verification and diarization, etc. However, most of these approaches are computationally intensive due to using transformer encoder and lack of sub-sampling. In this paper, we propose a new self-supervised learning model termed as Neural Encoder for Self-supervised Training (NEST). Specifically, we adopt the FastConformer architecture, which has an 8x sub-sampling rate and is faster than Transformer or Conformer architectures. Instead of clustering-based token generation, we resort to fixed random projection for its simplicity and effectiveness. We also propose a generalized noisy speech augmentation that teaches the model to disentangle the main speaker from noise or other speakers. Experiments show that the proposed NEST improves over existing self-supervised models on a variety of speech processing tasks. Code and checkpoints will be publicly available via NVIDIA NeMo toolkit [^1] https://github.com/NVIDIA/NeMo.
@@ I Introduction
Most recent speech self-supervised models are inspired by the BERT [1] model, which learn text token embedding by predicting the target of the masked positions given the context of the unmasked ones. Among them, are two main streams of contrastive and praedictive models. The contranstive approach [2, 3, 4, 5] quantizes the speech features into a set of target feature vectors and trains with a contrastive loss using the positive and negative target features. Meanwhile, the predictive approach [6, 7, 8, 9] quantizes the speech features into tokens and train with masked token prediction loss as in BERT [1]. In addition to the two approaches, some works also learn from the masked auto-encoding [10] approach and train speech self-supervised models with a reconstruction objective [11, 5].
@@
However, previous SSL models also have their different limitations. First, some models [2, 6, 8] use CNN+transformer architecture which has a short frame length of 25ms, which slows down the model’s inference speed. Second, the HuBERT-style quantization is very computationally expensive, which could take 20% of total training time according to XEUS [9]. Third, although BEST-RQ [7] uses Conformer [15] encoder with 40ms frame length and simple random quantization, it lacks the ability to explicitly tell one speaker from another, which limits its performance on speaker tasks (e.g., speaker diarization).
In this paper, we tackle all these challenges and bring the best practices from previous works, which constitute the proposed Neural Encoder for Self-supervised Training (NEST) framework. Our contributions are summarized as follows:
@@
  • Different from previous SSL works that focus mostly on downstream tasks with limited data, we also show that NEST can benefit speech recognition and translation even when data is sufficient.
@@ II-A Speech Encoder
Current SOTA speech SSL models [8, 6] mostly use transformer encoder [16] or Conformer [15] as speech encoder, which have either 25ms or 40ms frame length. Here we choose the more efficient FastConformer [17] which applies 8x convolutional sub-sampling on the input Mel-spectralgram before the following FastConformer layers, resulting in an 80ms frame length that can significantly reduce the sequence length to be processed by self-attention layers.
@@ II-C Speech Quantization
We resort to BEST-RQ [7] as our speech quantization method. Specifically, we use a single randomly initialized and frozen codebook of 8192 vocabulary and 16 dimension features. A randomly initialized and frozen linear layer is applied to the input Mel-spectralgram features to project them into the same dimension as the codebook, then a nearest neighbor search is applied to obtain the target tokens.
@@ II-D Feature Masking
We employ a random block-wise masking mechanism on the input Mel-spectralgram features, where each frame in the input has a probability $p_{m}$ as being selected as the start of a masking block. After randomly selecting a set of starting frames, we mask $l_{m}$ consecutive frames for each of the starting frames. Note that there could be overlapping between two masked blocks, which allows for arbitrary lengths in the resulting masked segments that do not overlap with each other. We use $p_{m}=0.01$ and $l_{m}=40$ in all our experiments.
@@ II-E Training
Since quantization is performed before the convolutional sub-sampling, there is a mismatch in length between the predicted tokens and target tokens. To mathc the sequence lengths, target tokens are averaged for every 8 frames, then apply threshold of 0.9 to select frames to be taken into loss calculation. Cross-entropy loss is applied on selected tokens determined by the input masks.
@@ III-B Results on SUPERB Multi-task Speech Processing
As presented in Table I, our NEST-L model is able to outperform WavLM-base++ [8] with similar size of parameters on all tasks, and also outperforms WavLM-large [8] that is 3x as large on speaker verification (SV), speaker diarization (SD) and phoneme recognition (PR). When compared with the XEUS [9] model that is trained on 10x data, we can see that our NEST-XL model is still able to achieve better performance on all speaker and content tasks, with especially large improvements on skeaper verification, speaker diarization and phoneme recognition. Overall, we are able to achieve new state-of-the-art results on SID, SV, SD, PR and ASR tasks compared with WavLM [8] that has data size as well as XEUS [9] that is trained on much large data, demonstrating the effectiveness of NEST when applied on various downstream speech processing tasks.
@@
Model Params Data (hrs) Europarl mExpresso Fleurs Average
@@
Model DIHARD3 CALLHOME-part2
@@ III-E Results on Speaker Diarization
To evaluate our model’s performance on speaker diarization, we follow the Sortformer [sortformer] architecture and attach 18 layers of transformer encoder [16] after the NEST encoder. We also train a simpler baseline with only FastConformer encoder and a linear decoder, with either random initialization or NEST initialization. For training data, we use a combination of real data (Fisher [24]) and simulated data (composed from LibriSeech [42], Fisher [24], Voxceleb1&2 [47], SRE [48, 49]) generated by the NeMo speech data simulator [50], with a total of 6.4K hours. We compare with state-of-the-art diarization models EEND-EDA [35] and EEND-GLA [34] and show the results in Table IV.
@@
SpeechBrain-Hubert-large LL-60K 89.37 80.54 77.44 78.96
ESPnet-Conformer N/A 86.30 N/A N/A 71.40

### v2

@@ Abstract
Self-supervised learning has been proved to benefit a wide range of speech processing tasks, such as speech recognition/translation, speaker verification and diarization, etc. However, most of these approaches are computationally intensive due to using transformer encoder and lack of sub-sampling. In this paper, we propose a new self-supervised learning model termed as pretraiNed Encoder for Speech Tasks (NEST). Specifically, we adopt the FastConformer architecture, which has an 8x sub-sampling rate and is faster than Transformer or Conformer architectures. Instead of clustering-based token generation, we resort to fixed random projection for its simplicity and effectiveness. We also propose a generalized noisy speech augmentation that teaches the model to disentangle the main speaker from noise or other speakers. Experiments show that the proposed NEST improves over existing self-supervised models on a variety of speech processing tasks. Code and checkpoints will be publicly available via NVIDIA NeMo toolkit [^1] https://github.com/NVIDIA/NeMo.
@@ I Introduction
Most recent speech self-supervised models are inspired by the BERT [1] model, which learn text token embedding by predicting the target of the masked positions given the context of the unmasked ones. Among them, are two main streams of contrastive and predictive models. The contrastive approach [2, 3, 4, 5] quantizes the speech features into a set of target feature vectors and trains with a contrastive loss using the positive and negative target features. Meanwhile, the predictive approach [6, 7, 8, 9] quantizes the speech features into tokens and train with masked token prediction loss as in BERT [1]. In addition to the two approaches, some works also learn from the masked auto-encoding [10] approach and train speech self-supervised models with a reconstruction objective [11, 5].
@@
However, previous SSL models also have their different limitations. First, some models [2, 6, 8] use CNN+transformer architecture which has a short frame length of 20ms, which slows down the model’s inference speed. Second, the HuBERT-style quantization is very computationally expensive, which could take 20% of total training time according to XEUS [9]. Third, although BEST-RQ [7] uses Conformer [15] encoder with 40ms frame length and simple random quantization, it lacks the ability to explicitly tell one speaker from another, which limits its performance on speaker tasks (e.g., speaker diarization).
In this paper, we tackle all these challenges and bring the best practices from previous works, which constitute the proposed pretraiNed Encoder for Speech Tasks (NEST) framework. Our contributions are summarized as follows:
@@
  • Different from previous SSL works that focus mostly on downstream tasks with limited data, we also show that NEST can benefit speech recognition and translation even when data is relatively larger.
@@ II-A Speech Encoder
Current SOTA speech SSL models [8, 6] mostly use transformer encoder [16] or Conformer [15] as speech encoder, which have either 20ms or 40ms frame length. Here we choose the more efficient FastConformer [17] which applies 8x convolutional sub-sampling on the input Mel-spectrogram before the following FastConformer layers, resulting in an 80ms frame length that can significantly reduce the sequence length to be processed by self-attention layers.
@@ II-C Speech Quantization
We resort to BEST-RQ [7] as our speech quantization method. Specifically, we use a single randomly initialized and frozen codebook of 8192 vocabulary and 16 dimension features. A randomly initialized and frozen linear layer is applied to the input Mel-spectrogram features to project them into the same dimension as the codebook, then a nearest neighbor search is applied to obtain the target tokens.
@@ II-D Feature Masking
We employ a random block-wise masking mechanism on the input Mel-spectrogram features, where each frame in the input has a probability $p_{m}$ as being selected as the start of a masking block. After randomly selecting a set of starting frames, we mask $l_{m}$ consecutive frames for each of the starting frames. Note that there could be overlapping between two masked blocks, which allows for arbitrary lengths in the resulting masked segments that do not overlap with each other. We use $p_{m}=0.01$ and $l_{m}=40$ in all our experiments.
@@ II-E Training
Since quantization is performed before the convolutional sub-sampling, there is a mismatch in the lengths between the predicted tokens and target tokens. To mathc the sequence lengths, target tokens are averaged for every 8 frames, then apply threshold of 0.9 to select frames to be taken into loss calculation. Cross-entropy loss is applied on selected tokens determined by the input masks.
@@ III-B Results on SUPERB Multi-task Speech Processing
As presented in Table I, our NEST-L model is able to outperform WavLM-base++ [8] with similar size of parameters on all tasks, and also outperforms WavLM-large [8] that is 3x as large on speaker verification (SV), speaker diarization (SD) and phoneme recognition (PR). When compared with the XEUS [9] model that is trained on 10x data, we can see that our NEST-XL model is still able to achieve better performance on all speaker and content tasks, with especially large improvements on speaker verification, speaker diarization and phoneme recognition. Overall, we are able to achieve new state-of-the-art results on SID, SV, SD, PR and ASR tasks compared with WavLM [8] that has data size as well as XEUS [9] that is trained on much large data, demonstrating the effectiveness of NEST when applied on various downstream speech processing tasks.
@@
Model DIHARD3 eval CALLHOME-part2
@@ III-E Results on Speaker Diarization
To evaluate our model’s performance on speaker diarization, we follow the Sortformer [sortformer] architecture and attach 18 layers of transformer encoder [16] after the NEST encoder. We also train a simpler baseline with only FastConformer encoder and a linear decoder, with either random initialization or NEST initialization. For training data, we use a combination of real data (Fisher [24]) and simulated data (composed from LibriSeech [42], SRE [47, 48]) generated by the NeMo speech data simulator [49], with a total of 6.4K hours. We compare with state-of-the-art diarization models EEND-EDA [35] and EEND-GLA [34] and show the results in Table IV.
@@
SpeechBrain-Hubert-large [51] LL-60K 89.37 80.54 77.44 78.96
ESPnet-Conformer [52] N/A 86.30 N/A N/A 71.40